Write the slope-intercept equation for the line that passes through (-8, -10) and is perpendicular to -8x – 3y = -16
I am stuck now. So far I have..
y-int: -16/-3= 5 1/3 (0, 5 1/3)
x-int: -16/-8= 2 (2,0)
Is this right so far?? Whats next? Slope?How do you plot it?
Thanks so MUCH!!!!

Answer by Alan3354(69443)

Write the slope-intercept equation for the line that passes through (-8,-10) and is perpendicular to -8x – 3y = -16
I am stuck now. So far I have..
y-int: -16/-3= 5 1/3 (0, 5 1/3)
x-int: -16/-8= 2 (2,0)
Is this right so far?? Whats next? Slope?How do you plot it?
==========
I don't know what you're doing.
-----------
Find the slope of -8x – 3y = -16 by solving for y:
y = (-8/3)x + something, doesn't matter
The slope is -8/3
The slope of lines perpendicular is the negative reciprocal, 3/8
------------
Use y = mx + b and the point to find b, the y-int
-10 = (3/8)*-8 + b
b = -7
--------
--> y = (3/8)x - 7 is the equation
============
There's no mention plotting in the problem.
=========================
PS Don't put a space after the comma in points (x,y)